2. Find the equation of the tangent line to the curve f(x)=x^{2}+5x at the point where x = 4. Show work
Answer #1

sohntion : Given the cower, cure: f(x) = x2 +5% Af x=G: Flo: 614) = 42 +544= 36. ] ..x=4 corresponds to the point (4,f(4)) =(if satisfied with the explanation, please rate it up..